OpenStudy (anonymous):

What is the solution to the following system? -2x-y=1 -4x-2y=-1

OpenStudy (anonymous):

set one equation equal to y, then substitute it into the other equation -2x-y=1 -4x-2y=-1 -4x-2y=-1 -2y = -1+4x y = (-1+4x)/-2 y = (-1/-2)+(4x/-2) y = (1/2) + -2x now plug this into the other untouched equation -2x-y=1 -2x-((1/2) + -2x)=1 now solve for x! once you get x, you can solve for y by plugging in x into either of the original equations.